Update - We are making progress!

AAPT/TPG fibre has been installed and equipment commissioned. Telstra is progressing with final transmission acceptance testing. Service record details have been reviewed and confirmed for migration. Telstra has determined the cable order of the 11 main cables to be migrated which is expected to commence later this week. AAPT/TPG NOC resources will monitor the ports as services are migrated. Customer services are expected to gradually restore as Telstra works through the main cable migration work. No firm ETR has been provided by Telstra for the completion of the main cable migrations, however we anticipate this to be relatively quick from this point on.
Aug 14, 12:56 AEST
Update - We has been advised that Telstra continues to progress with the rebuild of the Parklea exchange. There is no timeframe available yet for the completion of this work.

Telstra has completed the final Power installation work. AAPT/TPG fibre installation work is expected to be completed by 03/08/2017. The splicing of the fibre at the exchange has been delayed by Telstra due to copper cabling works and to avoid potential damage to the fibre cable. Telstra has advised that the connection of copper based services is expected to commence on 07/08/2017 and is dependent on the coordination with other carriers. No firm ETR has been provided by Telstra for service restoration.
Aug 3, 10:02 AEST
Update - We have been advised that Telstra continues to progress with the rebuild of the Parklea exchange. There is no timeframe available yet for the completion of this work.

Telstra has advised that the DC power system has been commissioned. Final mains power install is expected to be completed by the end of 21/07/2017 which will provide the power requirements for the exchange. Fibre installation work is ongoing with no ETR for completion available. Cable installation to the MDF has commenced and is ongoing. Copper cutovers have not commenced and no ETR has been provided for this work.

We will update with further information next week as it is provided to us, and thank our customers for their continued patience while this major piece of infrastructure is rebuilt.
Jul 14, 11:33 AEST
Update - AAPT/TPG has advised us that Telstra continues to progress with the rebuild of the Parklea exchange. There is no timeframe available yet for the completion of this work.

Telstra has advised that the temporary structures have been delivered to the Parklea site. Temporary entry and power is currently being progressed. The installation of equipment, internal cabling and power system works is expected to continue this week and possibly next. This will be followed by the installation of street cabling in preparation for the cutover of services onto the new exchange. No firm ETR has been provided by Telstra for service restoration.

The next update will be provided on 03/07/2017 or as soon as further information becomes available.
Jun 27, 12:17 AEST
Update - AAPT/TPG has been advised that Telstra continues to progress with the rebuild of the Parklea exchange. There is no timeframe available yet for the completion of this work.

Telstra has advised that the temporary structures have been scheduled to be relocated to the Parklea site on 22/06/2017. The integration of the structures to the exchange is expected to occur. The relevant drawings are then expected to be provided to AAPT/TPG to allow the Fibre and Field Technicians to plan and reinstate Fibre work. The 3rd party provider is expected to commence the re-termination of Copper cabling and complete the jumpering to the AAPT/TPG equipment over the following week. No ETR has been provided by Telstra for service restoration.

The next update will be provided on 26/06/2017 or as soon as further information becomes available.
Jun 20, 13:09 AEST
Update - AAPT/TPG has been advised that Telstra continues to progress with the rebuild of the Parklea exchange. There is no timeframe available yet for the completion of this work.

Telstra has authorised the final work to be completed on the temporary structures. This work involves the installation of the MDF and Power to the AAPT/TPG rack. AAPT/TPG technicians are scheduled to complete this work by 15/06/2017. The temporary structures are expected to be transported to the Parklea site by 16/06/2017. Integration of the structures to the exchange foundation is expected to commence with no ETR available from Telstra for this work as yet.

An ETR for total service restoration is unavailable. The next update will be provided on 19/06/2017 or as soon as further information becomes available.
Jun 13, 14:33 AEST
Update - AAPT/TPG has been advised that Telstra continues to progress with the rebuild of the Parklea exchange. There is no timeframe available yet for the completion of this work.

The site has been cleared and council approvals have been granted which has allowed the exchange foundation (concrete slab) to be completed. The work continues in preparation for the prefabricated structures to be delivered and installed. Final assembly of cabling and power is expected to occur after the prefabricated structures are delivered.

AAPT/TPG has completed the installation of equipment at the Erskine Park staging site. The delivery of the prefabricated structures to the Parklea site has not been confirmed but is possibly expected to occur by 09/06/2017.

An ETR for total service restoration is unavailable. The next update will be provided on 12/06/2017 or as soon as further information becomes available.
Jun 7, 12:21 AEST
Update - AAPT/TPG has been advised that Telstra continues to progress with the rebuild of the Parklea exchange. There is no timeframe available yet for the completion of this work.

While the rebuild continues at Parklea, equipment is being prepared at a temporary staging site in Erskine Park. AAPT/TPG staff have been approved for induction and installation of equipment into the Erskine Park site on 30/05/2017.

Final assembly of cable and power at the Parklea site is expected to occur when the replacement structures are delivered from the temporary staging site. No ETA has been provided on when the temporary structures are to be delivered to the Parklea site.

An ETR for total service restoration is unavailable. The next update will be provided on 05/06/2017 or as soon as further information becomes available.
May 29, 11:56 AEST
Update - AAPT have further advised today that they expect the rebuild of core infrastructure at the exchange to require a further two weeks, at which point AAPT can commence restoring services.
May 19, 09:29 AEST
Update - Telstra's service status has updated to indicate service restoration between Friday 19th May and Monday 29th May, however these dates are subject to change. AAPT have advised they expect to be restoring services next week, however this is dependent on site cleanup.
May 17, 09:48 AEST
Update - Telstra have advised that some services have been restored over the weekend, however we anticipate that issues relating to this outage will continue for several weeks due to the complexity of the event and the complete destruction of the exchange caused by the fire.
May 15, 10:18 AEST
Update - Telstra are still waiting for the arrival of the MEOW. We have also been advised of a substantial amount of toxic materials on site which need to be removed before service restoration can occur.
May 10, 14:03 AEST
Update - Telstra are deploying the Mobile Exchange on Wheels (MEOW) from Victoria and it should arrive today.
May 10, 11:07 AEST
Update - Some photos of the external damage: https://twitter.com/big_gaffa/status/861778669324382208
May 9, 14:16 AEST
Update - Site investigations have been completed, and Telstra has determined they are deploying a mobile exchange to the site to begin to restore services. There is no Estimated Time for Restoration at this stage.
May 9, 13:11 AEST
Identified - Telstra has advised of a fire in the Parklea Exchange overnight. The exchange is currently inoperable. Services fed from this exchange are offline at present with no ETR.
May 9, 09:10 AEST
Network Services Operational
90 days ago
100.0 % uptime
Today
IP Transit Operational
90 days ago
100.0 % uptime
Today
Ethernet services Operational
90 days ago
100.0 % uptime
Today
NBN Operational
90 days ago
100.0 % uptime
Today
ADSL Services Operational
90 days ago
100.0 % uptime
Today
Fixed Wireless Ethernet Operational
90 days ago
100.0 % uptime
Today
Intercapital Capacity Operational
90 days ago
100.0 % uptime
Today
3G/4G Data Services Operational
90 days ago
100.0 % uptime
Today
Private WAN (MPLS VPN) Operational
90 days ago
100.0 % uptime
Today
Managed VPN Operational
90 days ago
100.0 % uptime
Today
Voice Services Operational
90 days ago
100.0 % uptime
Today
Voice trunking Operational
90 days ago
100.0 % uptime
Today
Hosted voice Operational
90 days ago
100.0 % uptime
Today
ISDN Services Operational
Telstra Rebill Operational
Email Services Operational
90 days ago
100.0 % uptime
Today
SpamCloud Operational
90 days ago
100.0 % uptime
Today
Hosted Exchange Operational
90 days ago
100.0 % uptime
Today
POP/IMAP Legacy Email Operational
Hosting and Data Centre Operational
90 days ago
100.0 % uptime
Today
Virtual Machine Hosting Operational
90 days ago
100.0 % uptime
Today
Co-location Operational
90 days ago
100.0 % uptime
Today
Web Hosting Operational
90 days ago
100.0 % uptime
Today
Managed Backup (Dell Rapid Recovery) Operational
Managed Backup (Veeam) Operational
Billing Systems Operational
Helpdesk Services Operational
AWS cloudFront Operational
AWS ec2-ap-southeast-2 Operational
AWS ec2-ap-southeast-1 Operational
Xero Ltd - API Accounting API Operational
Xero Ltd - API Files API Operational
Xero Ltd - API Payroll API Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
System Metrics Month Week Day
Network Availability (syd01 Data Centre)
Fetching
Network Availability (syd02 Data Centre)
Fetching
Network Availability (mel01 Data Centre)
Fetching
Network Availability (cbr01 data centre)
Fetching
SIP Trunking Availability
Fetching
Hosted Voice Availability
Fetching
Past Incidents
Aug 19, 2017

No incidents reported today.

Aug 18, 2017

No incidents reported.

Aug 17, 2017

No incidents reported.

Aug 16, 2017
Resolved - Maintenance was completed at approx 1:30am
Aug 16, 07:26 AEST
Investigating - One of our upstream providers is currently performing network maintenance affecting some DSL and NBN services as a result of an earlier software bug in a routing platform. Estimated time for service restoration is 4am.
Aug 16, 00:06 AEST
Aug 15, 2017
Completed - The scheduled maintenance has been completed.
Aug 15, 04:00 AEST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Aug 14, 22:00 AEST
Update - Updated date: Due to an internal issue this work has been moved to the following Monday. Apologies about any confusion.
Jul 26, 18:02 AEST
Scheduled - In this window engineers will be migrating customer and internal gateways from legacy devices to our MX and ASR routers at NextDC S1. Please note while the outage window is large the outage per customer service should be minimal. For any questions regarding this work please email support@rwts.com.au.
Jul 26, 17:28 AEST
Aug 14, 2017
Resolved - This fault has been resolved.
Aug 14, 12:54 AEST
Monitoring - Our upstream provider has advised that the fault has been isolated to two devices within Optus' network. A workaround has been put in place and emergency maintenance has been completed on the first of these two as scheduled yesterday. We are advised that all calls are now flowing correctly and there should be no further impact as a result of this fault.

Further maintenance is scheduled for this evening at 23:00AEST on the second device to fully resolve the fault condition and restore full capacity and redundancy.
Aug 12, 11:37 AEST
Update - A failed carrier interconnect into the Optus network is causing the ongoing outage with inbound DID services in Victoria and the Northern Territory. Currently work is scheduled at 23:00AEST to completely resolve the outage by remediating the affected devices. We continue to apologise for the disruption.
Aug 11, 15:41 AEST
Identified - We have seen a reoccurrence of the issue with one way audio to Melbourne DIDs reported yesterday and confirmed resolved by Vocus again today. Vocus have acknowledged the fault and are continuing to work on this as a priority.
Aug 11, 09:26 AEST
Aug 13, 2017

No incidents reported.

Aug 11, 2017
Resolved - The complete fix has been implemented and we are now seeing correct audio flow.
Aug 11, 07:43 AEST
Monitoring - We have been advised that a fix was put in place at 3:45pm and all calls should now be working correctly.
Aug 10, 16:48 AEST
Update - This incident may also be impacting other state DID's. We are working with our provider on a restoration for these services
Aug 10, 14:22 AEST
Identified - One of our upstream providers have advised of a priority 1 incident causing one way audio to a range of potted and assigned DIDs. The outage originally occurred at 9am and was resolved at 9:15 but has reoccurred mid morning and remains ongoing.

Our provider has advised that they are currently experiencing a priority 1 network outage impacting multiple customers.

We have escalated this issue with our Service Manager to ensure that we are given priority for restoration.

While our provider is working to influence call routing around the affected circuits however, they are inbound calls and we are dependant on our provider repairing this issue.
Aug 10, 14:04 AEST
Aug 9, 2017

No incidents reported.

Aug 8, 2017

No incidents reported.

Aug 7, 2017

No incidents reported.

Aug 6, 2017

No incidents reported.

Aug 5, 2017

No incidents reported.